gelphielvr:

(algebra 1) slope intercept, y=mx+b does anyone know what formula I have to use to solve this? (question in the replies)

gelphielvr:

1 attachment
korbie:

do you have a graph with it, or are you solving it in standard slope intercept form?

gelphielvr:

@korbie wrote:
do you have a graph with it, or are you solving it in standard slope intercept form?
theres no graph involved just standard slope

korbie:

ok give my about 5 minutes

korbie:

y=-5/2x-3 I'm pretty sure

gelphielvr:

@korbie wrote:
y=-5/2x-3 I'm pretty sure
what formula did you use/how did you solve it? cause I'm totally lost on the process

korbie:

so if your tryna do perpendicular the easier trick is to flip the coefficient and donominator and add a negative so instead of 2/5 its -5/2 then you add one to the slope "+2" and make it negative

ThanosQC:

To solve this equation, you need to first find the negative reciprocal of the current "m" value. Since m = 2/5, the negative reciprocal of this would be -5/2. Now that you have what m=, you can now solve for b. To do this, plug in the x and y coordinates of the point you need to pass into the formula y=mx+b, in this case, (2, -2), and isolate. -2 = -5/2(2) + b -2+5 = -5+5 + b b = 3 Now you have all of the information you need to write the equation of the perpendicular line: m = -5/2 and b = 3; y = -5/2x + 3

Cypres:

trust me dude once you get the feeling of it and understand how simple it is and you don't let all the variables trick you its rlly simple, good luck tho